# ProxySPEX: Inference-Efficient Interpretability via Sparse Feature Interactions in LLMs

### v1

@@ 5.2 Model Component Attribution
We study the role of attention heads for a question-answering task using Llama-3.1-8B-Instruct and MMLU (high-school-us-history), which is a multiple-choice dataset. We treat each attention head as a feature and aim to identify interactions among heads using ProxySPEX. Let $L$ represent the number of layers in an LLM and let $\mathcal{L}\subseteq[L]$ represent a subset of the layers. Let $\mathcal{H}_{\mathcal{L}}$ denote the set of attention heads within these layers. For a subset of heads $S\subseteq\mathcal{H}_{\mathcal{L}}$, we set the output of heads in $\mathcal{H}_{\mathcal{L}}\setminus S$ to 00 and denote the ablated LLM as $\text{LLM}_{S}(\cdot)$. Define $f$ as:
@@
Pruning results. We use the LASSO and ProxySPEX to identify the most important heads for various sparsity levels ( i.e., the number of retained heads) across different sets of layers. We also compare to a Best-of-$N$ baseline, where we take the best of $N=5000$ different randomly chosen $S$, further details are in Section C.2. We use the procedure detailed in Section 4.2 to identify heads to remove for both ProxySPEX and LASSO. Test accuracies for each method are presented in Figure 9 at three different sparsity levels, and with three different layer ranges: initial ($1$-$3$), middle ($14$-$16$) and final ($30$-$32$). We observe that ProxySPEX consistently outperforms both baselines, with a higher test accuracy on the pruned models identified using ProxySPEX.
Characterizing interactions between attention heads. Analyzing the Fourier spectrum learned by ProxySPEX offers insights into the nature of the internal mechanisms of the LLM. As shown in Figure 9 (bottom), the spectral energy attributed to interactions, particularly within-layer interactions, markedly increases in deeper layers of Llama-3.1-8B-Instruct. There are many works that look at the differing functional roles of attention heads across layers [56]. ProxySPEX provides an exciting new quantitative approach to further investigate these phenomena.
Figure 9: Attention head pruning for Llama-3.1-8B-Instruct for MMLU (high-school-us-history). Top: We report the test accuracy vs. percentage of heads retained, comparing ProxySPEX, LASSO, and Best-of-$N$ across layer groups ($1$-$3$, $14$-$16$, $30$-$32$). Unpruned accuracy shown by dashed line.
Bottom: ProxySPEX’s learned spectral energy distribution into linear effects, within-layer, and across-layer interactions per layer group.
@@ 6 Discussion
Conclusion. We introduce ProxySPEX, an inference-efficient interaction attribution algorithm that efficiently scales with $n$ by leveraging an observed hierarchical structure among significant interactions in the Fourier spectrum of the model. Experiments across $4$ high-dimensional datasets show that ProxySPEX exploits hierarchical interactions via a GBT proxy model to reduce inferences by $\sim$10×10$\times$10 × over SPEX [4] while achieving equally faithful explanations. We demonstrate the importance of efficient interaction discovery by applying ProxySPEX to data and model component attribution.
Limitations. GBTs effectively capture hierarchical interactions but may not perform as well when interactions have a different structure. For example, simulations in Section B.4 empirically confirm that GBTs suffer in the case of sparse but non-hierarchical functions. More generally, in cases where the proxy GBT model is not faithful, the interactions identified by ProxySPEX might not be representative of the model’s reasoning. Another limitation is the degree of human interpretability that can be understood from computed interactions. While interactions can offer richer insights, they are more difficult to parse than marginal alternatives. Further improvements in visualization and post-processing of interactions are needed to fully harness the advances of ProxySPEX.
@@ A.1 Fourier Conversions
Banzhaf Interaction $I^{\text{B}}(T)$ $-2F(T)$
@@
Shapley Taylor $I^{\text{ST}}_{\ell}(T)$$\begin{cases}I^{\text{M}}(T),&|T|<\ell,\\ \vspace{10pt}\displaystyle\sum_{S\supseteq T}\binom{|S|}{\ell}^{-1}\,I^{M}(S),&|T|=\ell.\end{cases}$
@@
Faith-Shapley $I^{\text{FS}}_{\ell}(T)$ $I^{M}(T)+(-1)^{\ell-|T|}\frac{|T|}{\ell+|T|}\binom{\ell}{|T|}\sum_{\begin{subarray}{c}S\supset T\\ |S|>\ell\end{subarray}}F(S)\,\gamma(S,T,\ell)$ where $\text{where }\gamma(S,T,\ell)=\sum_{\begin{subarray}{c}T\subset R\subseteq S\\ |R|>\ell\end{subarray}}\frac{\binom{|R|-1}{\ell}}{\binom{|R|+\ell-1}{\ell+|T|}}(-2)^{|R|}$
@@ A.2 Fourier Extraction
17:              $v_{L}\leftarrow\mathcal{N}_{L}[S]$triangleright Mapping returns 0 if not contained
@@ A.3 Sparse Fourier Optimization
Equivalently, the function can be represented (and efficiently converted) under the Möbius transform. Converting Fourier to Möbius (via Section A.1), letting $\mathcal{K}^{+}\;=\;\bigl{\{}\,R\subseteq T\,\bigm{|}\,T\in\mathcal{K}\bigr{\}}$, and applying the inverse Möbius transform:
@@
The first constraint guarantees that whenever a monomial is activated (i.e. $x_{i}=1\hskip 5.0pt\forall i\in R$), all of its subsets are also activated. The second constraint ensures that if a monomial is deactivated (i.e. $\exists\hskip 2.0pti\in R\hskip 4.0pts.t.\hskip 4.0ptx_{i}=0$), at least one of its constituent terms ($y_{\{i\}}$) is likewise deactivated. After the optimization is solved, the solution can be read-off from the univariate monomials $y_{\{i\}}$. These monomial terms can also be used to impose cardinality constraints on the solution, as was used in Section 4.2 and Section 5.2.
@@ B.2 Measuring Spectral Hierarchies
Figure 10: (top row) We run SPEX until $R^{2}>0.9$. We report the faithfulness of when we truncate the spectrum to keep just the top $k$ coefficients for a range of $k$. We include results from Sentiment $n\in[256,512]$, and MS-COCO $n\in[60,85]$. In both cases faithfulness steadily increases as we increase $k$. (middle row) We report the $SCR$ (10) for the same top $k$ Fourier truncated functions above. In all cases, the $SCR$ is nearly $100\%$. (bottom row) We also report the $SHR$ (11), which is the strongest of the metric we consider. Here we find that even though $SHR$ decreases somewhat as $k$ grows, it is still strongly in favor of the hierarchy hypothesis.
@@
Figure 10 visualizes these rates alongside faithfulness ($R^{2}$) for the Sentiment Analysis and MS-COCO datasets. These empirical results aim to demonstrate that LLM feature interactions exhibit significant hierarchical structure. The high $SCR$ and $SHR$ scores support the core motivation for ProxySPEX: that important interactions are often built upon their lower-order subsets, a structure that Gradient Boosted Trees (GBTs) are well-suited to capture and exploit.
@@ B.3 Sparsification
The process of sparsification is crucial for enhancing the interpretability of the explanations generated by ProxySPEX . By retaining only the top $k$ Fourier coefficients, we can achieve a more concise and understandable representation of the model’s behavior without significantly compromising the faithfulness of the explanation. As demonstrated in Figure 5, a relatively small number of Fourier coefficients (approximately $200$) are often sufficient to achieve faithfulness comparable to using a much larger set of coefficients for tasks like sentiment classification and image captioning (MS-COCO).
Further results in Figure 11 illustrate the relationship between relative faithfulness and Fourier sparsity for both Sentiment and MS-COCO datasets across different inference multipliers ($\alpha$). These plots show that faithfulness generally increases with $k$, plateauing after a certain number of coefficients, reinforcing the idea that a sparse representation can effectively capture the essential dynamics of the LLM’s decision-making process.
@@ B.4 Proxy Model Selection
Figure 12 provides a comparative view of proxy model performance. Figure 12(a) and Figure 12(b) illustrate the faithfulness ($R^{2}$) of different proxy models (LASSO, Random Forest, Neural Network, and GBTs) on both a synthetic dataset with a complete hierarchy (defined below) and the Sentiment Analysis dataset, respectively, across various inference parameters ($\alpha$). These results empirically support the superiority of GBTs in capturing these complex interaction structures. However, it’s also important to acknowledge limitations; for instance, GBTs may not perform as well when interactions possess a different, non-hierarchical sparse structure, as empirically confirmed by simulations like the Synthetic-Peak example (which lacks hierarchical structure) shown in Figure 12(c).
@@
where $\mathcal{P}$ is a set of 10 uniformly where =$\mathcal{H}=\left\{R\subseteq T\,|\,T\in\mathcal{P}\right\}$
sampled sets of cardinality 5 and $F(R)\sim$ Uniform(1,1)11(-1,1)( - 1 , 1 ) for $R\in\mathcal{H}$
and $F(T)\sim$ Uniform(1,1)11(-1,1)( - 1 , 1 ) for $T\in\mathcal{P}$
@@
B.5 Practical Implications
@@
Figure 13 presents the practical benefits in terms of wall clock time for achieving different levels of faithfulness ($R^{2}$) on the Sentiment Analysis (Figure 13(a)) and MS-COCO (Figure 13(b)) datasets. These plots clearly demonstrate the speedups achieved by ProxySPEX. For example, in the sentiment analysis task using the smaller DistilBERT model, ProxySPEX offers a speedup of approximately 3x, while for the larger CLIP-ViT-B/32 model with MS-COCO, the speedup is around 5x when compared to methods that require more extensive sampling. This increased efficiency makes ProxySPEX a more viable tool for interpreting complex LLMs in real-world scenarios where computational resources and time are often constrained.
@@ C.1 Data Attribution via Non-Linear Datamodels
The training masks and margin outputs were provided by [52], corresponding to their subsampling rate of 50% (i.e., half the training images were used to fit each model). See [52] for the hyper-parameters selected. With $n=50{,}000$ training samples, $300{,}000$ training masks (model retrainings) were provided. This corresponds to $\alpha\approx 0.38$, which underscores the inference-efficiency of ProxySPEX  to identify strong interactions.
Utilizing these masks and margins, we randomly selected 60 test images (6 from each class) for analysis with ProxySPEX. Below, in Figure 14 and Figure 15, we present the strongest second-order interactions of the first thirty of these selected test images. Figure 8 visualizes the six test images exhibiting the most significant third-order interactions identified through this analysis.
After fitting ProxySPEX, we convert the Fourier interactions to Möbius using Section A.1. Since target and non-target images affect the test margin in opposite directions, we partition the interaction space into the following categories:
@@
$R^{*}$ =$=\begin{cases}S^{*}&\text{if }|I^{M}(S^{*})|\geq|I^{M}(T^{*})|\\ T^{*}&\text{otherwise}\end{cases}$
Visually, as presented in Figure 14 for $r=2$, the interactions $R^{*}$ identified by this rule often involve training images that appear to work together to reinforce or clarify the classification of the held-out image, frequently by contributing complementary features or attributes. It is important to acknowledge that this definition serves as a heuristic and does not perfectly isolate synergy; For example, the first frog image contains redundant bird images due to strong higher-order interactions involving these bird images.
@@
$R^{*}$ =$=\begin{cases}S^{*}&\text{if }|I^{M}(S^{*})|\geq|I^{M}(T^{*})|\\ T^{*}&\text{otherwise}\end{cases}$
Figure 15 demonstrates that this definition identifies redundant training images that are similar to the held-out image.
@@ C.2 Model Component Attribution
For an $L$ layer LLM, we let $[L]$ denote the set of layers and let $\mathcal{H}_{\ell}$ denote the set of attention heads in layer $\ell\in[L]$. For each experiment, we focus on a particular group of layers $\mathcal{L}\subseteq[L]$ within the model and denote the corresponding set of attention heads as $\mathcal{H}_{\mathcal{L}}=\bigcup_{\ell\in\mathcal{L}}\mathcal{H}_{\ell}$. The Llama-3.1-8B-Instruct model consists of $L=32$ layers, each with 32 attention heads.
@@
Pruning. We perform pruning experiments across three different layer groups $\mathcal{L}$: initial layers (=$\mathcal{L}=\{1,2,3\}$), middle layers (=$\mathcal{L}=\{14,15,16\}$), and final layers (=$\mathcal{L}=\{30,31,32\}$). Since each layer has 32 attention heads, we effectively perform ablation over $n=|\mathcal{H}_{\mathcal{L}}|=96$ features (attention heads) in total. For a given group $\mathcal{L}$, we begin by estimating the function $f_{\mathcal{L}}$ using both LASSO and ProxySPEX, based on evaluations of $f_{\mathcal{L}}(S)$ for $5000$ subsets $S$ sampled uniformly at random. These estimates serve as surrogates for the true head importance function. We then maximize the estimated functions to identify the most important attention heads under varying sparsity constraints (target numbers of retained heads). We use the procedure detailed in Section 4.2 to identify heads to remove for both ProxySPEX and LASSO. We also compare against a Best-of-$N$ baseline, in which the model is pruned by selecting the subset $S$ that achieves the highest value of $f_{\mathcal{L}}(S)$ among $5000$ randomly sampled subsets at the target sparsity level.
@@
In Figure 9, we report the value of $g_{\mathcal{L}}(S)$ for the pruned models obtained by each method. We find that ProxySPEX consistently outperforms both baselines, yielding higher test accuracy across all evaluated sparsity levels.

### v2

@@ 5.2 Model Component Attribution
We study the role of attention heads for a question-answering task using Llama-3.1-8B-Instruct and MMLU (high-school-us-history), which is a multiple-choice dataset. We treat each attention head as a feature and aim to identify interactions among heads using ProxySPEX. Let $L$ represent the number of layers in an LLM and let $\mathcal{L}\subseteq[L]$ represent a subset of the layers. Let $\mathcal{H}_{\mathcal{L}}$ denote the set of attention heads within these layers. For a subset of heads $S\subseteq\mathcal{H}_{\mathcal{L}}$, we set the output of heads in $\mathcal{H}_{\mathcal{L}}\setminus S$ to 0 and denote the ablated LLM as $\text{LLM}_{S}(\cdot)$. Define $f$ as:
@@
Pruning results. We use the LASSO and ProxySPEX to identify the most important heads for various sparsity levels ( i.e., the number of retained heads) across different sets of layers. We also compare to a Best-of-$N$ baseline, where we take the best of $N=5000$ different randomly chosen $S$, further details are in Section˜C.2. We use the procedure detailed in Section˜4.2 to identify heads to remove for both ProxySPEX and LASSO. Test accuracies for each method are presented in Figure˜10 at three different sparsity levels, and with three different layer ranges: initial (11-33), middle (1414-$16$) and final (3030-$32$). We observe that ProxySPEX consistently outperforms both baselines, with a higher test accuracy on the pruned models identified using ProxySPEX.
Characterizing interactions between attention heads. Analyzing the Fourier spectrum learned by ProxySPEX offers insights into the nature of the internal mechanisms of the LLM. As shown in Figure˜10 (bottom), the spectral energy attributed to interactions, particularly within-layer interactions, markedly increases in deeper layers of Llama-3.1-8B-Instruct. There are many works that look at the differing functional roles of attention heads across layers [59]. ProxySPEX provides an exciting new quantitative approach to further investigate these phenomena.
Figure 10: Attention head pruning for Llama-3.1-8B-Instruct for MMLU (high-school-us-history). Top: We report the test accuracy vs. percentage of heads retained, comparing ProxySPEX, LASSO, and Best-of-NN across layer groups (11-33, 1414-$16$, 3030-$32$). Unpruned accuracy shown by dashed line. Bottom: ProxySPEX’s learned spectral energy distribution into linear effects, within-layer, and across-layer interactions per layer group.
@@ 6 Discussion
Conclusion. We introduce ProxySPEX, an inference-efficient interaction attribution algorithm that efficiently scales with $n$ by leveraging an observed hierarchical structure among significant interactions in the Fourier spectrum of the model. Experiments across $4$ high-dimensional datasets show that ProxySPEX exploits hierarchical interactions via a GBT proxy model to reduce inferences by $\sim$$10\times$ over SPEX [4] while achieving equally faithful explanations. We demonstrate the importance of efficient interaction discovery by applying ProxySPEX to data and model component attribution.
Limitations. GBTs effectively capture hierarchical interactions but may not perform as well when interactions have a different structure. For example, simulations in Section˜B.4 empirically confirm that GBTs suffer in the case of sparse but non-hierarchical functions. More generally, in cases where the proxy GBT model is not faithful, the interactions identified by ProxySPEX might not be representative of the model’s reasoning. Another limitation is the degree of human interpretability that can be understood from computed interactions. While interactions can offer richer insights, they are more difficult to parse than marginal alternatives. Further improvements in visualization and post-processing of interactions are needed to fully harness the advances of ProxySPEX.
@@ A.1 Fourier Conversions
Banzhaf Interaction $I^{\text{B}}(T)$ 2F(T)-2F(T)
@@
Shapley Taylor $I^{\text{ST}}_{\ell}(T)$ $\begin{cases}I^{\text{M}}(T),&|T|<\ell,\\ \vskip 10.0pt\displaystyle\sum_{S\supseteq T}\binom{|S|}{\ell}^{-1}\,I^{M}(S),&|T|=\ell.\end{cases}$
@@
Faith-Shapley $I^{\text{FS}}_{\ell}(T)$
$I^{M}(T)+(-1)^{\ell-|T|}\frac{|T|}{\ell+|T|}\binom{\ell}{|T|}\sum_{\begin{subarray}{c}S\supset T\\ |S|>\ell\end{subarray}}F(S)\,\gamma(S,T,\ell)$
$\text{where }\gamma(S,T,\ell)=\sum_{\begin{subarray}{c}T\subset R\subseteq S\\ |R|>\ell\end{subarray}}\frac{\binom{|R|-1}{\ell}}{\binom{|R|+\ell-1}{\ell+|T|}}(-2)^{|R|}$
@@ A.2 Fourier Extraction
17:     $v_{L}\leftarrow\mathcal{N}_{L}[S]$ $\triangleright$ Mapping returns 0 if not contained
@@ A.3 Sparse Fourier Optimization
Equivalently, the function can be represented (and efficiently converted) under the Möbius transform. Converting Fourier to Möbius (via Section˜A.1), letting $\mathcal{K}^{+}\;=\;\bigl\{\,R\subseteq T\,\bigm|\,T\in\mathcal{K}\bigr\}$, and applying the inverse Möbius transform:
@@
The first constraint guarantees that whenever a monomial is activated (i.e. $x_{i}=1\hskip 5.0pt\forall i\in R$), all of its subsets are also activated. The second constraint ensures that if a monomial is deactivated (i.e. $\exists\hskip 2.0pti\in R\hskip 4.0pts.t.\hskip 4.0ptx_{i}=0$), at least one of its constituent terms ($y_{\{i\}}$) is likewise deactivated. After the optimization is solved, the solution can be read-off from the univariate monomials $y_{\{i\}}$. These monomial terms can also be used to impose cardinality constraints on the solution, as was used in Section˜4.2 and Section˜5.2.
@@ B.2 Measuring Spectral Hierarchies
Figure 11: (top row) We run SPEX until $R^{2}>0.9$. We report the faithfulness of when we truncate the spectrum to keep just the top $k$ coefficients for a range of $k$. We include results from Sentiment $n\in[256,512]$, and MS-COCO $n\in[60,85]$. In both cases faithfulness steadily increases as we increase $k$. (middle row) We report the SCRSCR (10) for the same top $k$ Fourier truncated functions above. In all cases, the SCRSCR is nearly $100\%$. (bottom row) We also report the SHRSHR (11), which is the strongest of the metric we consider. Here we find that even though SHRSHR decreases somewhat as $k$ grows, it is still strongly in favor of the hierarchy hypothesis.
@@
Figure˜11 visualizes these rates alongside faithfulness ($R^{2}$) for the Sentiment Analysis and MS-COCO datasets. These empirical results aim to demonstrate that LLM feature interactions exhibit significant hierarchical structure. The high $SCR$ and $SHR$ scores support the core motivation for ProxySPEX: that important interactions are often built upon their lower-order subsets, a structure that Gradient Boosted Trees (GBTs) are well-suited to capture and exploit.
@@ B.3 Sparsification
The process of sparsification is crucial for enhancing the interpretability of the explanations generated by ProxySPEX . By retaining only the top $k$ Fourier coefficients, we can achieve a more concise and understandable representation of the model’s behavior without significantly compromising the faithfulness of the explanation. As demonstrated in Figure˜5, a relatively small number of Fourier coefficients (approximately $200$) are often sufficient to achieve faithfulness comparable to using a much larger set of coefficients for tasks like sentiment classification and image captioning (MS-COCO).
Further results in Figure˜12 illustrate the relationship between relative faithfulness and Fourier sparsity for both Sentiment and MS-COCO datasets across different inference multipliers ($\alpha$). These plots show that faithfulness generally increases with $k$, plateauing after a certain number of coefficients, reinforcing the idea that a sparse representation can effectively capture the essential dynamics of the LLM’s decision-making process.
@@ B.4 Proxy Model Selection
Figure˜13 provides a comparative view of proxy model performance. Figure˜13(a) and Figure˜13(b) illustrate the faithfulness ($R^{2}$) of different proxy models (LASSO, Random Forest, Neural Network, and GBTs) on both a synthetic dataset with a complete hierarchy (defined below) and the Sentiment Analysis dataset, respectively, across various inference parameters ($\alpha$). These results empirically support the superiority of GBTs in capturing these complex interaction structures. However, it’s also important to acknowledge limitations; for instance, GBTs may not perform as well when interactions possess a different, non-hierarchical sparse structure, as empirically confirmed by simulations like the Synthetic-Peak example (which lacks hierarchical structure) shown in Figure˜13(c).
@@
where $\mathcal{P}$ is a set of 10 uniformly where $\mathcal{H}=\left\{R\subseteq T\,|\,T\in\mathcal{P}\right\}$
sampled sets of cardinality 5 and $F(R)\sim$ Uniform(1,1)(-1,1) for $R\in\mathcal{H}$
and $F(T)\sim$ Uniform(1,1)(-1,1) for $T\in\mathcal{P}$
@@
B.5 Shapley Value Approximation
We repeat the experiments of Section 4.3 under the metric of mean squared error relative to those computed by KernalSHAP under an inference budget of 10,000. In Figure 14, we find that ProxySPEX uniformly outperforms KernelSHAP within this tested range. Just as with recall, with large enough $\alpha$, KernelSHAP eventually surpasses ProxySPEX.
Figure 14: Shapley value mean square error after $\alpha\cdot n\log_{2}(n)$ inferences for multipliers $\alpha\in\{0.25,0.5,0.75,1.0\}$. Across all four tasks and multipliers, within the tested range, ProxySPEX provides a better approximation of the values computed under 10,000 inferences. Error bands indicate the standard deviation across ten different runs of the algorithms.
B.6 Practical Implications
@@
Figure˜15 presents the practical benefits in terms of wall clock time for achieving different levels of faithfulness ($R^{2}$) on the Sentiment Analysis (Figure˜15(a)) and MS-COCO (Figure˜15(b)) datasets. These plots clearly demonstrate the speedups achieved by ProxySPEX. For example, in the sentiment analysis task using the smaller DistilBERT model, ProxySPEX offers a speedup of approximately 3x, while for the larger CLIP-ViT-B/32 model with MS-COCO, the speedup is around 5x when compared to methods that require more extensive sampling. This increased efficiency makes ProxySPEX a more viable tool for interpreting complex LLMs in real-world scenarios where computational resources and time are often constrained.
@@ C.1 Data Attribution via Non-Linear Datamodels
The training masks and margin outputs were provided by [55], corresponding to their subsampling rate of 50% (i.e., half the training images were used to fit each model). See [55] for the hyper-parameters selected. With n=50,000n=50{,}000 training samples, 300,000300{,}000 training masks (model retrainings) were provided. This corresponds to $\alpha\approx 0.38$, which underscores the inference-efficiency of ProxySPEX  to identify strong interactions.
Utilizing these masks and margins, we randomly selected 60 test images (6 from each class) for analysis with ProxySPEX. Below, in Figure˜16 and Figure˜17, we present the strongest second-order interactions of the first thirty of these selected test images. Figure˜9 visualizes the six test images exhibiting the most significant third-order interactions identified through this analysis.
After fitting ProxySPEX, we convert the Fourier interactions to Möbius using Section˜A.1. Since target and non-target images affect the test margin in opposite directions, we partition the interaction space into the following categories:
@@
[…]
Visually, as presented in Figure˜16 for $r=2$, the interactions $R^{*}$ identified by this rule often involve training images that appear to work together to reinforce or clarify the classification of the held-out image, frequently by contributing complementary features or attributes. It is important to acknowledge that this definition serves as a heuristic and does not perfectly isolate synergy; For example, the first frog image contains redundant bird images due to strong higher-order interactions involving these bird images.
@@
[…]
Figure˜17 demonstrates that this definition identifies redundant training images that are similar to the held-out image.
@@ C.2 Model Component Attribution
For an $L$ layer LLM, we let $[L]$ denote the set of layers and let $\mathcal{H}_{\ell}$ denote the set of attention heads in layer $\ell\in[L]$. For each experiment, we focus on a particular group of layers $\mathcal{L}\subseteq[L]$ within the model and denote the corresponding set of attention heads as $\mathcal{H}_{\mathcal{L}}$$\bigcup_{\ell\in\mathcal{L}}\mathcal{H}_{\ell}$. The Llama-3.1-8B-Instruct model consists of $L=32$ layers, each with 32 attention heads.
@@
Pruning. We perform pruning experiments across three different layer groups $\mathcal{L}$: initial layers ($\mathcal{L}=\{1,2,3\}$), middle layers ($\mathcal{L}=\{14,15,16\}$), and final layers ($\mathcal{L}=\{30,31,32\}$). Since each layer has 32 attention heads, we effectively perform ablation over $n=|\mathcal{H}_{\mathcal{L}}|=96$ features (attention heads) in total. For a given group $\mathcal{L}$, we begin by estimating the function $f_{\mathcal{L}}$ using both LASSO and ProxySPEX, based on evaluations of $f_{\mathcal{L}}(S)$ for $5000$ subsets $S$ sampled uniformly at random. These estimates serve as surrogates for the true head importance function. We then maximize the estimated functions to identify the most important attention heads under varying sparsity constraints (target numbers of retained heads). We use the procedure detailed in Section˜4.2 to identify heads to remove for both ProxySPEX and LASSO. We also compare against a Best-of-$N$ baseline, in which the model is pruned by selecting the subset $S$ that achieves the highest value of $f_{\mathcal{L}}(S)$ among $5000$ randomly sampled subsets at the target sparsity level.
@@
In Figure˜10, we report the value of $g_{\mathcal{L}}(S)$ for the pruned models obtained by each method. We find that ProxySPEX consistently outperforms both baselines, yielding higher test accuracy across all evaluated sparsity levels.